OpenStudy (anonymous):

ok so I need to sketch the straight line determined by the points and find the slope intercept equation of the line. (1,1) (2,4) I got everything except the book says the answer is y=3x-2 and im getting y=3x+1?

OpenStudy (anonymous):

Okay so you found out the slope is 3.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Then you plugged in the point \((1,1)\):\[ (1) = 3(1) + b \]

OpenStudy (anonymous):

i didn't do that.. but im following..

OpenStudy (anonymous):

You solved for \(b\):\[ 1-3 = b\implies -2 = b \]Then put it into the original equation: \[ y = 3x-2 \]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

can u just tell me what form of the line is that again..? is that point-slope?

OpenStudy (anonymous):

I don't know what you mean.

OpenStudy (anonymous):

hmm, like. what is the origional equation?

OpenStudy (anonymous):

ok, where did you get 1−3=b⟹−2=b from?

OpenStudy (anonymous):

actually, i get it. But why not use point (2,4)?

OpenStudy (anonymous):

The \(\implies\) is just like saying "the equation becomes this"

OpenStudy (anonymous):

You can use any point you want, if it is on the line.

OpenStudy (anonymous):

oh ok. Yea, i see the other one would work as well. Thanks!